OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA = 2(lw + lh + wh), where l is length, w is width, and h is height. What is the surface area (in square meters) of a rectangular prism with length 5 m, width 12 m, and height 6 m?

OpenStudy (johnweldon1993):

Well we have the formula right there...just plug in the numbers \[\large SA = 2((5 \times 12) + (5 \times 6) + (12 \times 6))\] \[\large SA = \space ?\]